OpenStudy (harrisonshark):

Solve for c. ab - c = 9 -ab - 9 9 + ab ab - 9

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D

OpenStudy (harrisonshark):

theres no d but sure

OpenStudy (anonymous):

ab-9 a=3, and b=3, and c=0 (ab-c=9)3*3-0=9 (ab-9) 3*3-9=0

OpenStudy (harrisonshark):

thank you :D

OpenStudy (anonymous):

Does it make sense now.

OpenStudy (harrisonshark):

yes

OpenStudy (anonymous):

Your welcome